# The number of ends of big mapping class groups

### v1

@@ 4.2. Full subsurfaces
Recall from Section 2.7 the fixed set $\mathcal{R}=\left\{{f_{1},\dots,f_{N},c_{1},\dots,c_{M}}\right\}$ of representatives of the equivalence classes of maximal ends of $S$ used to define the translatable curve graph.
@@
The above lemma is reminiscent of Lemma 3.6, but crucially, it is not as powerful because it requires $\alpha$ and $\beta$ to be disjoint. Lastly, we define the notions of flux-right and flux-left in the exact same way as in the previous section, and the statements and proofs of Lemma 3.20 all still hold in the more general setting.
@@ 4.3. Hamming distance
Let $\alpha$ and $\beta$ be separating curves. Suppose $\alpha(\eta,\lambda)$ is a lasso curve such that $\eta$ lies on the same side of $\alpha$ and $\beta$, and $C_{\eta}$ has genus 0 if $S$ has finite genus. Then $H(\alpha(\eta,\lambda),\beta)=H(\alpha,\beta)+\left|{{\mathrm{End}}_{d}(C_{\eta})}\right|$.
@@ 4.4. Main theorem
If $\Sigma$ is a stable avenue surface with at least one discrete-type end, then $\operatorname{Map}(\Sigma)$ is one-ended.
@@
Like in the proof of Theorem 3.21, the action of $\operatorname{Map}(\Sigma)$ on $\Gamma$ allows us to assume that $\beta$ is the boundary curve of some copy of $S$. Then we may use the translation $h$ to push $\beta$ as far towards $e_{+}$ as we want. In total, we may assume without loss of generality that $\beta$ lies on the right side of $\alpha$ such that $\alpha,\beta]$ is full and $d(\beta,o)\geq 3R$. Then Lemma 4.5 gives a straight path $(\gamma_{0}=\alpha,\gamma_{1},\dots,\gamma_{k}=\beta)$ from $\alpha$ and $\beta$. Since $\Sigma$ has at least one class of discrete-type ends, there exists a lasso $(\eta,\lambda)$ such that $\eta\subset\beta_{+}\cap o_{+}$, $C_{\eta}$ is homeomorphic to $\widehat{T}_{i}$ for some $1\leq i\leq n$, and $\lambda$ intersects each $\gamma_{i}$ exactly once. Consider the lasso path $(\gamma_{0}(\eta,\lambda),\dots,\gamma_{k}(\eta,\lambda))$. Note that $\alpha$ and $\gamma_{0}(\eta,\lambda)$ are neighbors in $\Gamma$, and so are $\beta$ and $\gamma_{k}(\eta,\lambda)$. Also, for each $i=0,\dots,k$, Lemma 4.13 implies that $H(\gamma_{i}(\eta,\lambda),o)=H(\gamma_{i},o)+1$. Since the lasso path is again a straight path, this procedure may be iterated indefinitely, and each iteration produces a new lasso path whose vertices have a Hamming distance from $o$ increased by 1. Repeat the procedure for a total of $2R$ iterations and denote by $(\gamma_{0}^{\prime},\dots,\gamma_{k}^{\prime})$ the final path obtained after the last iteration. Then for each $i=0,\dots,k$,
@@
So $(\gamma_{0}^{\prime},\dots,\gamma_{k}^{\prime})$ is disjoint from $B$. Furthermore, the endpoints of the lasso paths constructed in each iteration forms a path from $\alpha$ to $\gamma_{0}^{\prime}$ and a path from $\beta$ to $\gamma_{k}^{\prime}$. The first $R+1$ vertices along these paths have a distance of at least $2R$ from $o$ because $d(\alpha,o),d(\beta,o)\geq 3R$. The remaining vertices along these paths have a distance greater than $R$ from $o$ because the Hamming distance is greater than $R$. So these two paths are also disjoint from $B$. The concatenation of all three paths then is a path from $\alpha$ to $\beta$ which is disjoint from $B$. ∎

### v2

@@ 4.2. Full subsurfaces
Recall from Section 2.7 the set $\mathcal{R}=\left\{{f_{1},\dots,f_{N},c_{1},\dots,c_{M}}\right\}$ of representatives of the equivalence classes of maximal ends of $S$ used to define the translatable curve graph of $\Sigma=S^{\natural\mathbb{Z}}$.
@@
The next lemma is reminiscent of Lemma 3.6, but crucially, it is not as powerful because it requires $\alpha$ and $\beta$ to be disjoint.
@@
Lastly, we define the notions of flux-right and flux-left in the exact same way as in the previous section, and the statements and proofs of Lemma 3.20 all still hold in the more general setting.
@@ 4.3. Hamming distance
Let $\alpha$ and $\beta$ be separating curves. Suppose $\alpha(\eta,\lambda)$ is a lasso curve such that $\eta$ lies on the same side of $\alpha$ and $\beta$, and $C_{\eta}$ has genus 0 if $S$ has finite genus. Then
$H(\alpha(\eta,\lambda),\beta)=H(\alpha,\beta)+\left|{{\mathrm{End}}_{d}(C_{\eta})}\right|.$
@@ 4.4. Main theorem
If $\Sigma$ is a stable avenue surface with a discrete-type end, then $\operatorname{Map}(\Sigma)$ is one-ended.
@@
Like in the proof of Theorem 3.21, the transitive action of $\operatorname{Map}(\Sigma)$ on $\Gamma$ allows us to assume that $\beta$ is the boundary curve of some copy of $S$. Then we may use the translation $h$ to push $\beta$ as far towards $e_{+}$ or $e_{-}$ as we want. In total, we may assume without loss of generality that $\beta$ lies on the right side of $\alpha$ such that $\alpha,\beta]$ is full and $d(\beta,o)\geq 3R$. Then Lemma 4.5 gives a straight path $(\gamma_{0}=\alpha,\gamma_{1},\dots,\gamma_{k}=\beta)$ from $\alpha$ and $\beta$. Since $\Sigma$ has at least one class of discrete-type ends, there exists a lasso $(\eta,\lambda)$ such that $\eta\subset\beta_{+}\cap o_{+}$, $C_{\eta}$ is homeomorphic to $\widehat{T}_{i}$ for some $1\leq i\leq n$, and $\lambda$ intersects each $\gamma_{i}$ exactly once. Consider the lasso path $(\gamma_{0}(\eta,\lambda),\dots,\gamma_{k}(\eta,\lambda))$. Note that $\alpha$ and $\gamma_{0}(\eta,\lambda)$ are neighbors in $\Gamma$, and so are $\beta$ and $\gamma_{k}(\eta,\lambda)$. Also, for each $i=0,\dots,k$, Lemma 4.13 implies that $H(\gamma_{i}(\eta,\lambda),o)=H(\gamma_{i},o)+1$. Since the lasso path is again a straight path, this procedure may be iterated indefinitely, and each iteration produces a new lasso path whose vertices have a Hamming distance from $o$ increased by 1. Repeat the procedure for a total of $2R$ iterations and denote by $(\gamma_{0}^{\prime},\dots,\gamma_{k}^{\prime})$ the final path obtained after the last iteration. Then for each $i=0,\dots,k$,
@@
So $(\gamma_{0}^{\prime},\dots,\gamma_{k}^{\prime})$ is disjoint from $B$. Furthermore, the endpoints of the lasso paths constructed in each iteration yield a path $P$ from $\alpha$ to $\gamma_{0}^{\prime}$ and a path $Q$ from $\beta$ to $\gamma_{k}^{\prime}$. For each vertex $u$ among the first $R+1$ vertices along $P$, $d(u,o)\geq 2R$ because $d(\alpha,o),d(\beta,o)\geq 3R$. For each vertex $v$ among the remaining vertices of $P$,
$d(v,o)\geq H(v,o)\geq H(\alpha,o)+R+1>R.$
So $P$, and similarly $Q$, are also disjoint from $B$. The concatenation of all three paths then is a path from $\alpha$ to $\beta$ which is disjoint from $B$. ∎